# Rayleigh-Bénard convective motion of stratified fluids in the Earth’s troposphere

### v1

@@ Abstract
Recently, Kaladze and Misra [Phys. Scr. 99 (2024) 085013] showed that the tropospheric stratified fluid flows may be unstable by the effects of the negative temperature gradient and the temperature-dependent density inhomogeneity arising from the thermal expansion. They also predicted that the modification in the Brunt-Väisälä frequency by the density inhomogeneity can lead to Rayleigh-Bénard convective instability in the tropospheric unbounded layers. The purpose of the present work is to revisit the Rayleigh-Bénard convective instability in more detail by considering both unbounded and bounded tropospheric layers. We show that the conditions for instability in these two cases significantly differ. The critical values of the Raleigh numbers and the expressions for the instability growth rates of thermal waves in the two cases are obtained and analyzed. In the case of the bounded region, we also derive the necessary boundary conditions and note that the vertical wave number is quantified, and the corresponding eigenvalue problem is well-set.
I Introduction
@@
Convective instability of atmospheric fluids has several important effects, including affecting the vertical motion of stratified fluids [1]. As the instability develops, the vertical movements of fluids tend to become larger due to thermal expansion, resulting in a large-scale turbulent fluid flow, the formation of extensive vertical clouds, or severe weather conditions such as thunderstorms. In this context, several attempts have been made to develop the relevant theories of incompressible atmospheric stratified fluids under gravity to study the instability conditions by the effects of thermal expansion, the density inhomogeneity, the background thermal gradient as well as the interplay between the buoyancy force and the dissipation due to kinematic viscosity and/or thermal diffusivity [2, 3, 4, 5]. In the subsequent developments of the theory of atmospheric fluid dynamics, Kaladze and Misra [6, 7] showed how the temperature-dependent density inhomogeneity due to thermal expansion and background thermal gradient can significantly modify the well-known Brunt-Väisälä (BV) frequency, associated with the internal gravity waves, and can play a significant role for the onset of instability of stratified fluids.
@@
Recently, Kaladze and Misra [7] studied the influence of temperature-dependent density inhomogeneity due to thermal expansion on the stability of atmospheric stratified fluids owing to its important applications in the dynamics of atmospheric waves and instability affecting Earth’s climate system, and showed significant modification in the Brunt-Väisälä (BV) frequency. As a result, the necessary conditions for the instability were modified from the previous investigation [6]. They also predicted the possibility of the emergence of Rayleigh-Bénard convective instability of incompressible flows in unbounded tropospheric layers.
@@
The purpose of the present work is to revisit the previous theory [7] and advance it by exploring the theory of Rayleigh-Bénard convective motion and instability in details in both unbounded and bounded domains of the troposphere. We show that the instability conditions and the critical values of the Rayleigh number in unbounded and bounded layers significantly differ. While the vertical wave number can assume any values in an unbounded domain, the same in a bounded domain is related to an eigenvalue problem. The corresponding instability growth rates are obtained and analyzed. We also show the existence of zero-frequency waves and neutral waves with a finite real wave frequency and zero growth or damping rate.
II Basic equations
@@
We consider the convective motion of incompressible stratified neutral gas under gravity in the Earth’s troposphere ($0-15$ km) with the effects of the vertical temperature gradient $(d\overline{T}/dz)$ and the temperature-dependent density inhomogeneity due to thermal expansion. Typically, all atmospheric fluid motions are subject to the thermal lapse rate, $\Gamma=-d\overline{T}/dz$. However, we distinguish convective waves from convective instability by assuming a real wave frequency for the former and a zero real wave frequency but a finite growth rate to mean the latter. Our starting point is the following set of fluid equations for the perturbed quantities in the Boussinesq approximation [7].
@@
where ${\bf r}=(x,y,z)$, and the term proportional to $\beta$ in Eq. (1) represents the buoyancy force, mainly responsible for the convective motion. Furthermore, in Eq. (3), the thermal lapse rate $\Gamma$ can be positive (negative) for a negative (positive) temperature gradient and $\chi$ is the coefficient of the thermal diffusivity. For some relevant details, readers are referred to the work of Kaladze and Misra [7].
Following Ref. [7] and assuming the length scale of variation of $\overline{\rho}(z)$ is much larger than that of $p^{\prime}$, we obtain the following evolution equation for the perturbed fluid pressure.
@@
where $N^{2}$ is the Brunt-Väisälä frequency, given by [7],
@@
Here, $\rho_{0}\equiv\rho_{0}(z)$ is the inhomogeneous fluid mass density at $\overline{T}(z)=T_{0}$ (where $T_{0}$ can be considered as a reference temperature relevant for internal gravity waves) and ${\cal T}\equiv{\cal T}(z)=\overline{T}(z)-T_{0}$. We note that the modification of the BV frequency occurs by the effects of thermal expansion (proportional to $\beta$.) Our aim is not to discuss more about the characteristics of the BV frequency already in Ref. [7] but to study its role on the convective wave motion and instability as in the following two subsections II.1 and II.2 by considering both unbounded and bounded layers in the troposphere. To this end, we assume that the length scale of fluid density inhomogeneity is approximately a constant, and so is the squared frequency $N^{2}$.
To mention, the convective instability in an unbounded domain can manifest in different ways. Although the norm of the perturbations grows in time, the perturbations can decay locally at every point in the unbounded domain, i.e., the growing perturbations are transported or convected towards infinity. Also, in many situations, including experiments and numerical simulations, although bounded domains are more relevant than unbounded ones, from a physical point of view, it is pertinent to investigate the instability conditions and the associated critical values of the Rayleigh number at different atmospheric conditions in an unbounded domain and to distinguish the relevant results with those in a bounded one.
II.1 Convective flow in an unbounded domain
To study the propagation characteristics of the convective wave motion, we assume that the pressure perturbation ($p^{\prime}$) propagates as a plane wave with the wave vector ${\bf k}=(k_{x},k_{y},k_{z})={\bf k}_{\perp}+k_{z}\hat{z}$ and the wave frequency $\omega$. Also, since the perturbations can be transported through an infinite extent, the wave amplitude $(p_{0})$ can be taken to be constant, i.e., independent of the vertical coordinate $z$. Thus, we consider
@@ II.1 Convective flow in an unbounded domain
Although constant amplitude is reasonably a good assumption for an unbounded domain, it can, however, vary with $z$ in the bounded domain, to be discussed later in Sec. II.2. Substituting Eq. (7) into Eq. (5), we obtain the following dispersion relation for Rayleigh-Bénard convective waves [7].
@@
In what follows, contributions of different physical parameters and forces can be noted. From Eq. (8), it is clear that the thermal diffusivity (proportional to $\chi$) and the kinematic viscosity (proportional to $\nu$) effects give rise to wave damping in absence of the buoyancy force (proportional to $\beta$). On the other hand, without these dissipative effects, the convective wave motion can be unstable by the effects of the modified BV frequency $N^{2}$ [7]. Before proceeding to study the general dispersion relation (See Case IV) for the characteristics of convective modes, we first consider some particular cases of interest (See Cases I to III). This will enable us to identify different important contributions from different physical parameters, and to distinguish them with the influence of $N^{2}$.
@@
In a most simple situation, i.e., in absence of the dissipative (thermal diffusivity and viscosity) effects and when the background density $\overline{\rho}$ is independent of the vertical coordinate $z$, i.e., $N^{2}=0$, the dispersion equation (8) reduces to
@@
Clearly, the propagating mode with real wave frequency exists in the case when $\Gamma<0$ with positive thermal gradient, i.e., in the stratospheric region [6]. We can call such waves without any instability growth rate as convective modes. On the other hand, when $\Gamma>0$, which holds for a negative thermal gradient, such as for stratified fluids in the troposphere (See Table 1 of Ref. [6]), the convective motion can be unstable with the instability growth rate, given by,
@@
where we have considered only the positive sign in the right-hand side because, since there is no dissipative effects due to the thermal diffusion or kinematic viscosity, the wave can not be damped. From Eq. (10), it is also evident that the convective instability is mainly caused by the buoyancy force, arising due to thermal expansion of vertically stratified fluids under gravity with a negative thermal gradient (or, $\Gamma>0$). Furthermore, the instability growth rate typically depends on the vertical wave number $k_{z}$ (Since $\gamma$ varies inversely with $k_{z}$ hidden in $k$), i.e., the convective flows become more unstable with an enhanced growth rate as the vertical scale or the wavelength of perturbations tend to increase.
@@
We move on to the case when the background density may not be constant but varies with $z$ such that $N^{2}\neq 0$ and without any dissipative effects. This case is similar to one previously studied by Kaladze and Misra [7] [See Eqs. (50) and (51) therein]. However, for limpidness, we repeat it here in our discussion. In this case, the BV frequency will specifically play a key role in determining the instability growth in the presence of the Buoyancy force. Thus, Eq. (8) reduces to
@@
Here, $K_{N}=\left(k^{4}+N^{4}k_{z}^{2}/g^{2}\right)^{1/4}$. Although the forms of $\gamma$ presented here and in Ref. [7] look different, they agree after using Eq. (14) and a minor correction with the factor $1/2$, which is missing in Eq. (52) of Ref. [7]. We note that Eqs. (12) and (13) are valid when $\omega_{r}\gamma<0$, i.e., when $\omega_{r}$ and $\gamma$ have the opposite signs. Thus, for a positive growth rate, $\gamma>0$ and $\omega_{r}<0$. The latter may be indicative of a phase shift in the perturbation or that the phase decreases with time. It is also important to note that, the growth rate is independent of the sign of $N^{2}$, implying that the convective flow without any dissipation in an unbounded domain may become unstable irrespective of the sign of the vertical background density gradient positive or negative. From Eqs. (12) and (13), we find that the perturbations due to convective fluid motion can propagate with a finite real wave frequency (unless at $k_{z}=0$, where the frequency vanishes) but are unstable having a finite growth rate $\gamma>0$. Interestingly, such a growth rate of convective waves appears to be larger than the frequency $\omega_{r}$ (Compare the square root factors in $\omega_{r}$ and $\gamma$). The larger value of the growth rate can be associated with the aperiodic fluid motion and higher values of the Rayleigh number (to be defined shortly) where dominant effects of the buoyancy force over the viscous force come into the picture. For a typical tropospheric model [6], we have $|N^{2}|\sim 10^{-4}-10^{-3}/\rm{s}^{2}$ and $g\sim 10~{}\rm{m/s^{2}}$ such that $K_{N}\approx k$ for $k_{z}\ll 1$. In this situation, $\omega_{r}\approx 0$, and one can have the dominant convective instability. Thus, in absence of the fluid viscosity and thermal diffusivity, the Brunt-Väisälä frequency, associated with the length scale of the background density inhomogeneity and modified by the temperature-dependent density inhomogeneity due to thermal expansion, can lead to the Rayleigh-Bénard convective instability with a finite growth rate $\gamma~{}(>0)$, but no cut-off at a finite value of $k$. From Eq. (13), we also find that as the vertical wave number $k_{z}$ becomes smaller or the corresponding vertical wavelength of perturbations tend to increase, the instability growth rate increases. Such an argument is in contrast to the instability condition to be studied in Case III below by defining the Rayleigh number. Furthermore, the instability growth rate increases with increasing values of the thermal expansion coefficient $\beta$ exceeding a critical value $\beta_{c}\sim 0.004$/K [7], which corresponds to $N^{2}<0$, i.e., the occurrence of the instability of stratified fluids in the troposphere [7, 6]. Thus, it may be concluded that the unstable fluid flow in tropospheric stratified layers associated with internal gravity waves may correspond to the convective instability due to the key roles of the buoyancy force in which the growing perturbations (with time) are transported towards infinite boundaries.
@@
From Eq. (16), we find that while the upper sign gives a purely damped mode due to dissipation $(\gamma\equiv\gamma_{-})$, the lower sign can give a buoyancy-force (dominating over the dissipation) driven convective instability with $\gamma\equiv\gamma_{+}>0$ when the second term (square root term) becomes larger in magnitude than the first term [proportional to $(\chi+\nu)$] in the square brackets, i.e., when
@@
Such a critical wave number typically depends on the ratio of the contributions from the buoyancy force and the dissipative force. In particular, the convective waves, which are neither damped nor amplified (i.e., $\gamma=0$) can be called neutral waves. Thus, it is pertinent to introduce the Rayleigh number, which is the ratio between the contributions from the buoyancy force (providing the destabilizing effect) and the dissipation due to the kinematic viscosity and the thermal diffusivity (providing the stabilizing effect), defined by,
@@
where $H$ is the height of the tropospheric fluid layer and $\Gamma=-d\overline{T}/dz\sim\Delta T/H>0$. When the Rayleigh number is below its critical value, the fluid convection does not occur and the heat is transferred only through the thermal conduction. However, the convection starts at the critical number and the heat is transferred through convection when the Rayleigh number is above the critical value. Rayleigh showed that the convective instability can occur when the thermal gradient $\Delta T$ is large enough to exceed $Ra$ a certain value. In order to estimate such a value, we recast Eq. (17) as [8]
@@
To note, we have considered the vertical domain as unbounded, and it may be reasonable to consider $H$ as the vertical length-scale of perturbations. Thus, with the vertical wave number $k_{z}$, the expression on right-hand side of Eq. (20) assumes the smallest value at $k_{\perp}=k_{z}/\sqrt{2}$, and we have the critical (minimum) value of the Rayleigh number, above which the convective instability occurs, as
@@
We observe that the value of $Ra_{\min}$ gets significantly reduced in the limit of $k_{z}H\ll 1$, i.e., as the vertical wavelength of perturbation significantly exceeds the vertical length scale. This may be true when the vertical domain is unbounded and the wave number $k_{z}$ can assume any small values for which $Ra$ becomes smaller with smaller values of $k_{z}$ relative to the inverse of the vertical scale. In this case, the buoyancy force may not be so strong over the dissipation for which the convective instability may become prominent. Thus, the convective motions of stratified viscous fluids (with constant background density) under gravity in an unbounded domain become more unstable as the vertical scale $H$ of perturbations deepens compared to the wavelength of perturbation. However, we do not have any information about the relation between the components $k_{x}$ and $k_{y}$ of the wave number $k$ except a value of $k_{\perp}=\sqrt{k_{x}^{2}+k_{y}^{2}}$. Comparing the growth rate [Eq. (16)] with that obtained in Case I [Eq. (10)], we find that the growth rate gets reduced by the effects of the kinematic viscosity and thermal diffusivity, and the reduction can be significant at higher values of $\chi$ and $\nu$. Also, similar to Case II, the growth rate can be increased with increasing values of $\beta$. Physically, while the viscous force (proportional to $\nu$) and the thermal diffusion (proportional to $\chi$) have stabilizing roles, the buoyancy force (proportional to $\beta$) plays the destabilizing roles in the wave motion. So, as the values of $\chi$ and $\nu$ are reduced or a value of $\beta$ is increased, the wave tends to become more unstable with a higher growth rate of perturbations.

### v2

@@ Abstract
Recently, Kaladze and Misra [Phys. Scr. 99 (2024) 085013] showed that the tropospheric stratified fluid flows may be unstable by the effects of the negative temperature gradient and the temperature-dependent density inhomogeneity arising from the thermal expansion. They also predicted that the modification in the Brunt-Väisälä frequency by the density inhomogeneity can lead to Rayleigh-Bénard convective instability in the tropospheric unbounded layers. The purpose of the present work is to revisit the Rayleigh-Bénard convective instability in more detail by considering both unbounded and bounded tropospheric layers. Starting from a set of fluid equations for incompressible neutral fluids with temperature gradients and using the Boussinesq approximation, we derive the general dispersion relations for Rayleigh-Bénard convective waves in unbounded and bounded tropospheric domains and analyze them with some particular cases both analytically and numerically. We show that the conditions for instability in these two cases significantly differ. Furthermore, we obtain and analyze the critical values of the Raleigh numbers and the expressions for the instability growth rates of thermal waves in the two cases. In the case of the bounded region, we also derive the necessary boundary conditions and note that the vertical wave number is quantified, and the corresponding eigenvalue problem is well-set.
keywords:
[…]
1 Introduction
@@
Convective instability of atmospheric fluids has several key consequences, including affecting the vertical motion of stratified fluids (Chitre and Gokhale, 1973). As the instability develops, the vertical movements of fluids tend to increase due to thermal expansion, resulting in a large-scale turbulent fluid flow, the formation of extensive vertical clouds, or severe weather conditions such as thunderstorms. In this context, there were several attempts to develop the relevant theories of incompressible atmospheric stratified fluids under gravity to study the instability conditions by the effects of thermal expansion, the density inhomogeneity, the background thermal gradient as well as the interplay between the buoyancy force and the dissipation due to kinematic viscosity and/or thermal diffusivity (Zhang et al., 2024; Shmerlin and Kalashnik, 2013; Kameyama and Kinoshita, 2013; Kopp et al., 2021). In the subsequent developments of the theory of atmospheric fluid dynamics, Kaladze and Misra (Kaladze and Misra, 2023, 2024) showed how the temperature-dependent density inhomogeneity due to thermal expansion and background thermal gradient can significantly modify the well-known Brunt-Väisälä (BV) frequency, associated with the internal gravity waves, and can play a significant role for the onset of instability of stratified fluids. On the other hand, several authors have also studied and advanced the theory of Darcy-Bénard convective instability in porous media (See, e.g., (Turkyilmazoglu and Siddiqui, 2023a, b; Turkyilmazoglu and Duraihem, 2024)
@@
Recently, Kaladze and Misra (Kaladze and Misra, 2024) studied the influence of temperature-dependent density inhomogeneity due to thermal expansion on the stability of atmospheric stratified fluids owing to its significant applications in the dynamics of atmospheric waves and instability affecting Earth’s climate system and showed considerable modification in the Brunt-Väisälä (BV) frequency. As a result, the necessary conditions for the instability were modified from the previous investigation (Kaladze and Misra, 2023). They also predicted the possibility of the emergence of Rayleigh-Bénard convective instability of incompressible flows in unbounded tropospheric layers. It should be noted that the Rayleigh-Bénard convection can be represented as a Nambu-metriplectic problem by incorporating dissipative effects (Bihlo, 2008).
@@
The purpose of the present work is to revisit the previous theory (Kaladze and Misra, 2024) and advance it by exploring the theory of Rayleigh-Bénard convective motion and instability in details in both unbounded and bounded domains of the troposphere. We show that the instability conditions and the critical values of the Rayleigh number in unbounded and bounded layers significantly differ. While the vertical wave number can assume any values in an unbounded domain, the same in a bounded domain is an eigenvalue problem. The corresponding instability growth rates are obtained and analyzed. We also show the generation of zero-frequency waves and neutral waves with a finite real wave frequency and zero growth or damping rate.
2 Basic equations
@@
We consider the convective motion of incompressible stratified neutral gas under gravity in the Earth’s troposphere ($0-15$ km) with the effects of the vertical temperature gradient $(d\overline{T}/dz)$ and the temperature-dependent density inhomogeneity due to thermal expansion. Generally, all atmospheric fluid motions are subject to the thermal lapse rate, $\Gamma=-d\overline{T}/dz$. However, we distinguish convective waves from convective instability by assuming a real-wave frequency for the former and a zero real-wave frequency but a finite growth rate to mean the latter. Our starting point is the following set of fluid equations for the perturbed quantities in the Boussinesq approximation (Kaladze and Misra, 2024).
@@
where ${\bf r}=(x,y,z)$, and the term proportional to $\beta$ in Eq. (1) represents the buoyancy force, mainly responsible for the convective motion. Furthermore, in Eq. (3), the thermal lapse rate $\Gamma$ can be positive (negative) for a negative (positive) temperature gradient and $\chi$ is the coefficient of the thermal diffusivity. For some relevant details, readers are referred to the work of Kaladze and Misra [See Eqs. (12), (13), and (17) of (Kaladze and Misra, 2024)].
Following Ref. (Kaladze and Misra, 2024) and assuming the length scale of variation of $\overline{\rho}(z)$ is much larger than that of $p^{\prime}$, we obtain the following evolution equation for the perturbed fluid pressure.
@@
where $N^{2}$ is the Brunt-Väisälä frequency, given by (Kaladze and Misra, 2024),
@@
Here, $\rho_{0}\equiv\rho_{0}(z)$ is the inhomogeneous fluid mass density at $\overline{T}(z)=T_{0}$ (where $T_{0}$ can be considered as a reference temperature relevant for internal gravity waves) and ${\cal T}\equiv{\cal T}(z)=\overline{T}(z)-T_{0}$. We note that the modification of the BV frequency occurs by the effects of thermal expansion (proportional to $\beta$.) Our aim is not to discuss more about the characteristics of the BV frequency already in Ref. (Kaladze and Misra, 2024) but to study its role in the convective wave motion and instability as in the following two subsections 2.1 and 2.2 by considering both unbounded and bounded layers in the troposphere. To this end, we assume that the length scale of fluid density inhomogeneity is approximately constant, and so is the squared frequency $N^{2}$.
We mention that the convective instability in an unbounded domain can manifest in several ways. Although the norm of the perturbations grows in time, the perturbations can decay locally at every point in the unbounded domain, i.e., the growing perturbations are transported or convected towards infinity. Also, in many situations, including experiments and numerical simulations, although bounded domains are more relevant than unbounded ones, from a physical point of view, it is pertinent to investigate the instability conditions and the associated critical values of the Rayleigh number at different atmospheric conditions in an unbounded domain and to distinguish the relevant results with those in a bounded one.
2.1 Convective flow in an unbounded domain
To study the propagation characteristics of the convective wave motion, we assume that the pressure perturbation ($p^{\prime}$) propagates as a plane wave with the wave vector ${\bf k}=(k_{x},k_{y},k_{z})={\bf k}_{\perp}+k_{z}\hat{z}$ and the wave frequency $\omega$. Also, since the perturbations can have an infinite extent, the wave amplitude $(p_{0})$ can be taken as constant, i.e., independent of the vertical coordinate $z$. Thus, we consider
@@ 2.1 Convective flow in an unbounded domain
Although constant amplitude is reasonably a good assumption for an unbounded domain, it can vary with $z$ in the bounded domain, to be discussed later in Sec. 2.2. Substituting Eq. (7) into Eq. (5), we obtain the following dispersion relation for Rayleigh-Bénard convective waves (Kaladze and Misra, 2024).
@@
In what follows, we note the contributions of different physical parameters and forces. From Eq. (8), it is clear that the thermal diffusivity (proportional to $\chi$) and the kinematic viscosity (proportional to $\nu$) effects give rise to wave damping in the absence of the buoyancy force (proportional to $\beta$). On the other hand, without these dissipative effects, the convective wave motion can be unstable by the influences of the modified BV frequency $N^{2}$ (Kaladze and Misra, 2024). Before analyzing the general dispersion relation (See Case IV) for the characteristics of convective modes, we first consider some particular cases of interest (See Cases I to III). These will enable us to identify several significant contributions from different physical parameters and to distinguish them with the influence of $N^{2}$.
@@
In a most simple situation, i.e., in the absence of the dissipative (thermal diffusivity and viscosity) effects and when the background density $\overline{\rho}$ is independent of the vertical coordinate $z$, i.e., $N^{2}=0$, the dispersion equation (8) reduces to
@@
Clearly, the propagating mode with real wave frequency exists when $\Gamma<0$ with positive thermal gradient, i.e., in the stratospheric region (Kaladze and Misra, 2023). We can call such waves without any instability growth rate as convective modes. On the other hand, when $\Gamma>0$, which holds for a negative thermal gradient, such as for stratified fluids in the troposphere (See Table 1 of Ref. (Kaladze and Misra, 2023)), the convective motion can be unstable with the instability growth rate, given by,
@@
where we have considered only the positive sign on the right-hand side because there are no dissipative effects due to the thermal diffusion or kinematic viscosity, so the wave can not be damped. From Eq. (10), it is also evident that the convective instability is mainly caused by the buoyancy force, arising due to the thermal expansion of vertically stratified fluids under gravity with a negative thermal gradient (or, $\Gamma>0$). Furthermore, the instability growth rate typically depends on the vertical wave number $k_{z}$ (Since $\gamma$ varies inversely with $k_{z}$ hidden in $k$), i.e., the convective flows become more unstable with an enhanced growth rate as the vertical scale, or the wavelength of perturbations tend to increase.
@@
We move on to the case when the background density may not be constant but varies with $z$ such that $N^{2}\neq 0$ and without any dissipative effects. This case is similar to one previously studied by Kaladze and Misra (Kaladze and Misra, 2024) [See Eqs. (50) and (51) therein]. However, for limpidness, we repeat it here in our discussion. In this case, the BV frequency will specifically play a key role in determining the instability growth in the presence of the Buoyancy force. Thus, Eq. (8) reduces to
@@
Here, $K_{N}=\left(k^{4}+N^{4}k_{z}^{2}/g^{2}\right)^{1/4}$. Although the forms of $\gamma$ presented here and in Ref. (Kaladze and Misra, 2024) look different, they agree after using Eq. (14) and a minor correction with the factor $1/2$, which is missing in Eq. (52) of Ref. Kaladze and Misra (2024). We note that Eqs. (12) and (13) are valid when $\omega_{r}\gamma<0$, i.e., when $\omega_{r}$ and $\gamma$ have the opposite signs. Thus, for a positive growth rate, $\gamma>0$ and $\omega_{r}<0$. The latter may indicate a phase shift in the perturbation or that the phase decreases with time. It is also important to note that the growth rate is independent of the sign of $N^{2}$, implying that the convective flow without any dissipation in an unbounded domain may become unstable irrespective of the sign of the vertical background density gradient, positive or negative. From Eqs. (12) and (13), we find that the perturbations due to convective fluid motion can propagate with a finite real wave frequency (unless at $k_{z}=0$, where the frequency vanishes) but are unstable having a finite growth rate $\gamma>0$. Interestingly, such a growth rate of convective waves appears larger than the frequency $\omega_{r}$ (Compare the square root factors in $\omega_{r}$ and $\gamma$). The large value of the growth rate can be related to the aperiodic fluid motion and higher values of the Rayleigh number (to be defined shortly), where dominant effects of the buoyancy force over the viscous force come into the picture. For a typical tropospheric model (Kaladze and Misra, 2023), we have $|N^{2}|\sim 10^{-4}-10^{-3}/\rm{s}^{2}$ and $g\sim 10~{}\rm{m/s^{2}}$ such that $K_{N}\approx k$ for $k_{z}\ll 1$. In this situation, $\omega_{r}\approx 0$, and one can have the dominant convective instability. Thus, in absence of the fluid viscosity and thermal diffusivity, the Brunt-Väisälä frequency, associated with the length scale of the background density inhomogeneity and modified by the temperature-dependent density inhomogeneity due to thermal expansion, can lead to the Rayleigh-Bénard convective instability with a finite growth rate $\gamma~{}(>0)$, but no cut-off at a finite value of $k$. From Eq. (13), we also find that as the vertical wave number $k_{z}$ becomes smaller or the corresponding vertical wavelengths of perturbations tend to increase, the instability growth rate increases. Such an argument contradicts the instability condition, to be studied in Case III below, by defining the Rayleigh number. Furthermore, the instability growth rate increases with increasing values of the thermal expansion coefficient $\beta$ exceeding a critical value $\beta_{c}\sim 0.004$/K (Kaladze and Misra, 2024), which corresponds to $N^{2}<0$, i.e., the occurrence of the instability of stratified fluids in the troposphere (Kaladze and Misra, 2024, 2023). Thus, we may conclude that the unstable fluid flow in tropospheric stratified layers associated with internal gravity waves may correspond to convective instability due to the key roles of the buoyancy force in which the growing perturbations (with time) get transported towards infinite boundaries.
@@
From Eq. (16), we find that while the upper sign gives a purely damped mode due to the dissipation $(\gamma\equiv\gamma_{-})$, the lower sign can give a buoyancy-force (dominating over the dissipation) driven convective instability with $\gamma\equiv\gamma_{+}>0$ when the second term (square root term) becomes larger in magnitude than the first term [proportional to $(\chi+\nu)$] in the square brackets, i.e., when
@@
Such a critical wave number typically depends on the ratio of the contributions from the buoyancy and dissipative forces. In particular, the convective waves, which are neither damped nor amplified (i.e., $\gamma=0$), can be called neutral waves. Thus, it is pertinent to introduce the Rayleigh number, which is the ratio between the contributions from the buoyancy force (providing the destabilizing effect) and the dissipation due to the kinematic viscosity and the thermal diffusivity (providing the stabilizing effect), defined by,
@@
where $H$ is the height of the tropospheric fluid layer and $\Gamma=-d\overline{T}/dz\sim\Delta T/H>0$. When the Rayleigh number is below its critical value, the fluid convection does not occur, and the heat is transferred only through thermal conduction. However, the convection starts at the critical value, and the heat is transferred through convection when the Rayleigh number is above the critical value. Rayleigh showed that convective instability can occur when the thermal gradient $\Delta T$ is large enough to exceed $Ra$ a particular value. In order to estimate such a value, we recast Eq. (17) as (Satoh, 2004)
@@
We have considered the vertical domain unbounded, and it may be reasonable to consider $H$ as the vertical length scale of perturbations. Thus, with the vertical component of the wave number $k_{z}$, the expression on the right-hand side of Eq. (20) assumes the smallest value at $k_{\perp}=k_{z}/\sqrt{2}$, and we have the critical (minimum) value of the Rayleigh number, above which the convective instability occurs, as
@@
We observe that the value of $Ra_{\min}$ gets significantly reduced in the limit of $k_{z}H\ll 1$, i.e., as the vertical wavelength of perturbation significantly exceeds the vertical length scale. It may be true when the vertical domain is unbounded and the wave number $k_{z}$ can assume any small values for which $Ra$ becomes smaller with smaller values of $k_{z}$ relative to the inverse of the vertical scale. In this case, the buoyancy force may not be so strong over the dissipation for which the convective instability may become prominent. Thus, the convective motions of stratified viscous fluids (with constant background density) under gravity in an unbounded domain become more unstable as the vertical scale $H$ of perturbations deepens compared to the perturbation wavelength. However, we do not have any information about the relation between the components $k_{x}$ and $k_{y}$ of the wave number $k$ except a value of $k_{\perp}=\sqrt{k_{x}^{2}+k_{y}^{2}}$. Comparing the growth rate [Eq. (16)] with that obtained in Case I [Eq. (10)], we find that the growth rate gets reduced by the effects of the kinematic viscosity and thermal diffusivity, and the reduction can be significant at higher values of $\chi$ and $\nu$. Also, similar to Case II, the growth rate can be increased with increasing values of $\beta$. Physically, while the viscous force (proportional to $\nu$) and the thermal diffusion (proportional to $\chi$) have stabilizing roles, the buoyancy force (proportional to $\beta$) plays destabilizing roles in the wave motion. So, as the values of $\chi$ and $\nu$ get reduced or a value of $\beta$ increases, the wave tends to become more unstable with a higher growth rate of perturbations.